OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of (f + g)(−2)? a.−9 b.−6 c.−3 d.9

OpenStudy (across):

I already told you what to do.

OpenStudy (paxpolaris):

is that f of (g of x)?

OpenStudy (anonymous):

there for it would be 9

OpenStudy (lollylau):

never knew you can add up functions like that.

OpenStudy (paxpolaris):

NO, f of g(x) is written as \[(f ○ g)(x)=[f \left( g \left( x \right) \right)\] \[\left( f+g \right)(x)=f(x)+g(x)\]

OpenStudy (lollylau):

Oh.